# Learn to Refuse: Making Large Language Models More Controllable and Reliable through Knowledge Scope Limitation and Refusal Mechanism

### v2

@@ 3.5 Refusal Mechanism
The refusal mechanism in L2R judges whether a question $Q$ can be answered or not and refuses to answer if it deems the question unanswerable. Two types of refusal mechanisms in L2R work together to make this decision: soft refusal and hard refusal. The former is executed by the LLM itself, while the latter is set by humans and can be adjusted based on different situations.
@@
$I^{\text{hard}}_{i}=\min(C\cdot S)<\alpha,$ (2)
@@ Appendix A Experiment Settings
We choose gpt-3.5-turbo-0613 as the underlying large language model for L2R in all tests. The temperature is set to 0 to reduce instability, and $top\_p$ is set to 1 by default. The only hyperparameter in L2R, $\alpha$, which represents the threshold for hard refusal, is set to 0.75 by default. For llama2, we select the model version of Llama-2-70b-chat-hf.
Retrieval augmentation plays a crucial role in our L2R system. Initially, we use all-mpnet-base-v2 from the Sentence-BERT family Reimers and Gurevych (2019) to obtain embeddings for all knowledge texts. We select to employ L2 Euclidean distance to measure the similarity score between the question and candidate knowledge. The system retrieve the top $k$ related knowledge for a single query, with the default value of $k$ set to 4. Specifically, we employed FAISS (Facebook AI Similarity Search) Johnson et al. (2019) to efficiently retrieve related documents from a large-scale knowledge base.
The prompts for all LLMs used in L2R can be found in Appendix F.

### v3

@@ 3.5 Refusal Mechanism
The refusal mechanism in L2R judges whether a question $Q$ can be answered or not and refuses to answer if it deems the question unanswerable. Two types of refusal mechanisms in L2R work together to make this decision: soft refusal and hard refusal. The former is from LLM’s generation output and is executed by the LLM itself, while the latter is set by humans and can be adjusted based on different situations. We categorize refusals as “Soft” and “Hard” from a system perspective. A soft refusal is defined as one originating directly from the LLM, and it is variable and adjustable based on different LLMs and their prompts. In contrast, a hard refusal involves a backup method, which requires additional computation and comparison with a system-defined threshold. We consider this hard refusal significant because if the knowledge base support is insufficient, the L2R system will refuse to answer to avoid hallucination, regardless of the LLM’s perspective.
@@
$I^{\text{hard}}=\min_{1\leq j\leq k}\left(\dfrac{S_{j}}{C_{j}}\right)<\alpha$ (2)
@@ Appendix A Experiment Settings
We choose gpt-3.5-turbo-0613 as the underlying large language model for L2R in all tests. The temperature is set to 0 to reduce instability, and $top\_p$ is set to 1 by default. The hyperparameter $\alpha$, which represents the threshold for hard refusal, is set to 0.75 by default to simplify experiments. For llama2, we select the model version of Llama-2-70b-chat-hf.
Retrieval augmentation plays a crucial role in our L2R system. Initially, we use all-mpnet-base-v2 from the Sentence-BERT family Reimers and Gurevych (2019) to obtain embeddings for all knowledge texts. We select to employ L2 Euclidean distance to measure the similarity score between the question and candidate knowledge. The system retrieve the top $k$ related knowledge for a single query, with the default value of $k$ set to 4. Specifically, we employed FAISS (Facebook AI Similarity Search) Johnson et al. (2019) to efficiently retrieve related documents from a large-scale knowledge base. All the knowledge base is mined from the same LLM used later to answer questions.
@@
The prompts for all LLMs used in L2R can be found in Appendix F.
Appendix B Ablation Study
@@ Appendix D Hyperparameter Analysis: Threshold Selection in Hard Refusal
Figure 6: The Precision Recall Curve under the change of $\alpha$.
@@
We also generate the precision-recall curve based on different $\alpha$ in Figure 6. Precision is the same as accuracy in the experiments, which is the corrected answer over the number of provided answers. Recall is calculated by the corrected answer over the number of all questions. From an overall view of the figure, we can observe a trade-off between precision and recall with the change of $\alpha$. We can focus on the curve between the recall of 40 and 60, where it is more stable, and better for selecting an $\alpha$ value within this range.